# A multiverse model in dS wedge holography

### v2

@@ 3.5 Central dogma and quantum cosmology interpretation
The previous calculation shows a non-trivial island between the UV and IR branes for a single double-sided UV brane. A natural possibility is to have an island that crosses multiple patches, as shown in Fig. 5 (bottom). In that case, the von Neumann entropy is the $NS^{\rm aft\leavevmode\nobreak\ PT}$ found above, with $N$ the number of patches. However, the dominating saddle is the one where the island only intercepts the IR brane a single time, see Fig 5 (top), which is given by (65). In contrast, the island saddle encompassing different universes would allow for overlapping HRT surfaces once we introduce another observer on a different double-sided UV brane, shown also in Fig. 5 (bottom). Then, the theory might be in tension with bulk reconstruction, as there would be causally disconnected UV brane observers sharing a common entanglement wedge, corresponding to the bulk AdS space, as this would modify the commutation relations between the observables accessible to the different UV branes. This could lead to a violation of the no-cloning theorem in quantum mechanics as two observers could reconstruct the same information simultaneously.
@@
Figure 5: Island saddles for the multiverse model, where $\mathbf{R}_{1}$ (blue dot) and $\mathbf{R}_{2}$ (orange dot) represent different codimension-one bath regions in the UV brane at a given global time $\tau_{c}$ when the observers are collecting information in each brane universe. Top: Islands ($\mathbf{I}_{1}$, $\mathbf{I}_{2}$, blue and orange solid curves with the same color scheme as the corresponding bath regions) connecting a single side of the UV brane to an IR brane, and producing the dominant contribution to the entropy (65). Bottom: A different possibility where $I_{1}$ and $I_{2}$ always overlap.
We can view the island transitions in our model through the lenses of quantum cosmology, which shares several similarities to Aguilar-Gutierrez:2021bns ; Levine:2022wos  [^15] A technical difference with these other models is that our multiverse configuration has been prepared by gluing different pair of branes from the Hartle-Hawking state, such that one does not need to account for conical singularities.. The information available to a single universe where gravity is non-dynamical (i.e. the UV brane) and includes the observers at other universes due to the gluing conditions requiring that all the universes are completely symmetric.. The microscopic degrees of freedom can be probed by observers on the UV brane. Yet, the UV observers can only probe the regions available through the islands due to the coarse-graining of information in dS wedge holography, which prohibits communication between observers at different branes.
Thus, dS wedge holography provides a conceptually new insight into this problem in the sense that the coarse-graining of information in quantum cosmology, at least for the braneworld multiverse models, involves entanglement between the UV and IR energy modes. It would be of interest to see whether this observation occurs in an explicit higher dimensional model of false vacuum decay in eternal inflation, such as Garriga:2006hw .
@@ 4 Discussion
In this work, we have constructed a multiverse toy model based on the dS wedge holography proposal in Aguilar-Gutierrez:2023tic . The multiverse arises by joining, through the Israel junction conditions, an arbitrary number of dS2 branes cutting off the interior and exterior of AdS3 space. To have an interesting evolution that resembles false vacuum eternal inflation, while keeping the model exactly solvable, we “dressed” the IR brane as a near Nariai black hole using the dS JT gravity as the intrinsic gravity theory on it. In principle, we could add JT couplings to the UV brane as well, but as we showed this would not qualitatively change our study in any way.
We studied the coarse-graining of local measurements expected in false vacuum eternal inflation Aguilar-Gutierrez:2021bns within our toy model, by evaluating the fine-grained von Neumann entropy (and islands) with the HRT formula. Double holography allowed us to determine the entanglement entropy of Hawking radiation collected by a UV brane observer, which we associated with this coarse-graining, introduced by the presence of the ETW branes. The appearance of islands connecting the UV and IR modes also allows for a new perspective on this problem.
The presence of dS JT gravity on the IR brane modifies the Page curve transition and allows for information recovery with respect to a given dS braneworld observer. Interestingly, we find an avatar of a previous proposal Cohen:1998zx , where a generic effective field theory in curved spacetime is expected to involve a relation between UV and IR cutoffs, in order for consistency relations in the effective field theory description. With our model, one arrives at a similar conclusion, as the Page time depends on both UV and IR quantities. Finally, the resulting entanglement of Hawking modes captured by a UV brane observer also shows agreement with previous arguments about the central dogma for cosmological horizons and consistency with the no-cloning theorem Levine:2022wos . Namely, there are no overlapping island saddles when we allow for multiple Dirichlet brane observers.
@@
Firstly, most of our work on entanglement in the multiverse model has been focused on the AdS3 bulk perspective, and the experience of the UV brane observers. However, it would be interesting to have an interpretation of the coarse-graining in quantum cosmology from the time-like separated codimension-two defects on the interception between the branes (see the brane interceptions at $\tau=\pm\frac{\pi}{2}$ in Fig. 1), as this might provide a dS/CFT holographic perspective Strominger:2001pn for the coarse-graining of the UV and IR degrees of freedom.

### v3

@@ 3.5 Central dogma and quantum cosmology interpretation
The previous calculation shows a non-trivial island between the UV and IR region from the perspective of a single observer. A natural possibility is to have an island that crosses multiple patches, as shown in Fig. 5 (bottom). In that case, the von Neumann entropy is the $NS^{\rm aft\leavevmode\nobreak\ PT}$ found above, with $N$ the number of patches. However, the dominating saddle is the one where the island only intercepts the IR brane at a single time, see Fig 5 (top), which is given by (59). In contrast, the island saddle encompassing different universes would allow for overlapping HRT surfaces once we introduce another observer on a different double-sided UV region, shown also in Fig. 5 (bottom). Then, the theory might be in tension with bulk reconstruction, as there would be causally disconnected UV observers sharing a common entanglement wedge, corresponding to the bulk AdS space, as this would modify the commutation relations between the observables accessible to the different UV regions. This could lead to a violation of the no-cloning theorem in quantum mechanics as two observers could reconstruct the same information simultaneously.
@@
Figure 5: Island saddles for the multiverse model, where $\mathbf{R}_{1}$ (blue dot) and $\mathbf{R}_{2}$ (orange dot) represent different codimension-one bath regions in the finite cutoff UV at a given global time $\tau_{c}$ when the observers are collecting information in each brane universe. Top: Islands ($\mathbf{I}_{1}$, $\mathbf{I}_{2}$, blue and orange solid curves with the same color scheme as the corresponding bath regions) connecting a single side of the UV region to an IR brane, and producing the dominant contribution to the entropy (59). Bottom: A different possibility where $I_{1}$ and $I_{2}$ always overlap.
We can view the island transitions in our model through the lenses of quantum cosmology, which shares several similarities to Aguilar-Gutierrez:2021bns ; Levine:2022wos  [^17] A technical difference with these other models is that our multiverse configuration does not have conical singularities.. The information available to a single universe where gravity is non-dynamical (i.e. the finite UV cutoff region) and includes the observers at other universes due to the gluing conditions requiring that all the universes are completely symmetric.. The microscopic degrees of freedom can be probed by observers on the UV cutoff region. Yet, the observers can only probe the regions available through the islands due to the coarse-graining of information in dS wedge holography, which prohibits communication between observers at different branes.
Thus, the proposed extension of dS wedge holography provides a conceptually new insight into this problem in the sense that the coarse-graining of information in quantum cosmology, at least within our setting, involves entanglement between the UV and IR energy modes. It would be of interest to see whether this observation occurs in an explicit higher dimensional model of false vacuum decay in eternal inflation, such as Garriga:2006hw .
@@ 4 Discussion
In this work, we have constructed a multiverse toy model based on the dS wedge holography proposal in Aguilar-Gutierrez:2023tic , and $T^{2}$ deformations Hartman:2018tkw . The multiverse arises by generating several copies of global AdS3 space with a finite radial cutoff in the UV arising from turning a $T^{2}$ deformation in the dual theory, and a dS2 ETW brane cutting off the interior, which are glued together in a periodic matter through the Israel junction conditions. To have an interesting evolution that resembles false vacuum eternal inflation, while keeping the model exactly solvable, we “dressed” the IR brane as a near Nariai black hole using the dS JT gravity as the intrinsic gravity theory on it. In principle, we could add JT couplings to the finite UV cutoff region as well (after performing the dual deformation), but as we showed this would not qualitatively change our study in any way.
We studied the coarse-graining of local measurements expected in false vacuum eternal inflation Aguilar-Gutierrez:2021bns within our toy model, by evaluating the fine-grained von Neumann entropy (and islands) with the HRT formula. Double holography allowed us to determine the entanglement entropy of Hawking radiation collected by an observer in the UV cutoff, which we associated with this coarse-graining.
The presence of dS JT gravity on the IR brane modifies the Page curve transition and allows for information recovery with respect to a given dS braneworld observer. Interestingly, we find an avatar of a previous proposal Cohen:1998zx , where a generic effective field theory in curved spacetime is expected to involve a relation between UV and IR cutoffs, in order to satisfy consistency relations in the effective field theory description. With our model, one arrives at a similar conclusion, as the Page time depends on both UV and IR quantities. Furthermore, the interplay between UV and IR degrees of freedom is clear from the the RT surfaces connecting these regions, giving a non-vanishing entanglement entropy; the existence of the IR brane is what allows the appearance of the connected RT surfaces and the Page transition. Finally, the resulting entanglement of Hawking modes captured by a UV cutoff observer also shows agreement with previous arguments about the central dogma for cosmological horizons and consistency with the no-cloning theorem Levine:2022wos . Namely, there are no overlapping island saddles when we allow for multiple Dirichlet brane observers.
@@
Firstly, most of our work on entanglement in the multiverse model has been focused on the AdS3 bulk perspective, and the experience of the UV observers. However, it would be interesting to have an interpretation of the coarse-graining in quantum cosmology from the time-like separated codimension-two defects on the interception between the branes (see the brane interceptions at $\tau=\pm\frac{\pi}{2}$ in Fig. 1), as this might provide a dS/CFT holographic perspective Strominger:2001pn for the coarse-graining of the UV and IR degrees of freedom.